given 1/u+1/v=1/f with f as a constant. if f=10 cm and u decrease with the rate of 2 cm/second, find the rate of v when u=40 cm


1
Expert's answer
2021-11-22T15:11:30-0500

Given


"\dfrac{1}{u}+\dfrac{1}{v}=\dfrac{1}{f}"

Differentiate both sides with respect to "t"


"(\dfrac{1}{u}+\dfrac{1}{v})'=(\dfrac{1}{f})'"

"-\dfrac{u'}{u^2}-\dfrac{v'}{v^2}=0"

Solve for "v'"


"v'=-\dfrac{v^2}{u^2}u'"

"v=\dfrac{uf}{u-f}"

Substitute


"v'=-\dfrac{f^2}{(u-f)^2}u'"

Given "f=10\ cm, u'=-2\ cm/s, u=40\ cm"


"v'=-\dfrac{(10\ cm)^2}{(40\ cm-10\ cm)^2}(-2\ cm/s)=\dfrac{2}{9}\ cm/ s"



"v" increase with the rate of "\dfrac{2}{9}" cm/second.